# Generative Diffusion Model Driven Massive Random Access in Massive MIMO Systems

### v2

@@ VI-A4 Comparative Schemes
For AUD, the proposed method is compared with two baseline algorithms: AMP [13] and OMP [12]. Additionally, we also assess the performance of the VPL-AUDNet scheme trained with a fixed pilot length. For CE and DD, the following algorithms are included in the comparisons:
@@
$\bullet$ Perfect-CSI & ZF-DD: The CSI $\bm{H}_{a}$ is perfectly known, and employ ZF for DD;
$\bullet$ Perfect-CSI & OAMP-DD: The CSI $\bm{H}_{a}$ is perfectly known, and utilize OAMP [21] for DD;
$\bullet$ Pilot LS-CE & ZF-DD: Utilize pilots to perform LS CE, and use the estimated channel for ZF DD;
$\bullet$ Pilot LS-CE & OAMP-DD: Utilize pilots to perform LS CE, and use the estimated channel for OAMP [21] DD;
$\bullet$ Iter LMMSE-CE & OAMP-DD: Employ LMMSE for CE and OAMP for DD, with CE and DD iteratively executed to enhance the performance of each other [21];
$\bullet$ Iter Langevin CE & DD: Conduct iterative CE and DD using ALD [33];
@@
$\bullet$ Perfect-CSI & SDE-DD: The CSI $\bm{H}_{a}$ is known, and apply SDE solver for DD;
$\bullet$ Our Iter SDE CE & DD: The proposed Algorithm 2.
@@ VI-C Performance of JCEDD
Figure 10: NMSE and BER performance with respect to the received SNR for different CE and DD algorithms under 3GPP channels.
@@
Fig. 10 illustrates the performance curves of CE and DD under varying SNRs for different schemes. In the simulations, each user is allocated a pilot length of $L_{p}\!=\!15$ and a data length of $L_{d}\!=\!50$. As shown in Fig. 10(a), the iterative CE and DD schemes (i.e., Iter Langevin CE & DD, Iter LMMSE-CE & OAMP-DD, and Our Iter SDE CE & DD) achieve an NMSE gain of approximately 7 and 10 dB when compared to the non-iterative Pilot LS-CE & ZF-DD, and Pilot LS-CE & OAMP-DD schemes. This indicates a significant gain in CE performance by using detected data to refine CE. As SNR increases, Iter LMMSE-CE & OAMP-DD’s performance approaches that of LS-CE & Perfect-Data. Similarly, the performance of Our Iter SDE CE & DD (Algorithm 2) matches SDE-CE & Perfect-Data (Algorithm 1) for $\text{SNR}\!\geq\!8$ dB. When the data is not perfectly known, Our Iter SDE CE & DD outperforms Iter LMMSE-CE & OAMP-DD because it more effectively utilizes the prior distribution of the channel through a learned prior score function, whereas Iter LMMSE-CE & OAMP-DD relies only on the channel’s second-order statistical information via LMMSE. Additionally, Our Iter SDE CE & DD implements a PC sampler to enhance SDE solution performance beyond what Iter Langevin CE & DD algorithm can achieve with only a corrector.
@@
Fig. 10(b) compares the BER performance of various algorithms. Similar to Fig. 10(a), the non-iterative Pilot LS-CE & ZF-DD and Pilot LS-CE & OAMP-DD algorithms perform worse than the iterative algorithms. Our Iter SDE CE & DD scheme achieves performance close to that of Perfect-CSI & SDE-DD and surpasses the Iter LMMSE-CE & OAMP-DD and Iter Langevin CE & DD algorithms. At a BER of $10^{-4}$, Our Iter SDE CE & DD offers a 3 dB SNR gain over Iter LMMSE-CE & OAMP-DD. Furthermore, Perfect-CSI & SDE-DD outperforms Perfect-CSI & OAMP-DD and Perfect-CSI & ZF-DD algorithms. This is likely because the channel matrix elements for the 3GPP scenario do not meet the i.i.d. distribution assumptions for DD tasks, causing OAMP algorithms to underperform. On the other hand, the ZF algorithm does not exploit the prior distribution information of the constellation. In contrast, Perfect-CSI & SDE-DD thoroughly leverages the prior distribution of the data and the likelihood information of the received signals, resulting in superior performance.
@@
Figure 11: NMSE and BER performance with respect to the received SNR for different CE and DD algorithms under the MIMO Rayleigh channel.
@@
In Fig. 11, we conducted simulations using the MIMO Rayleigh channel model, where the channel prior is assumed to follow $\mathcal{N}(\bm{0},\bm{I})$. Instead of utilizing a neural network-learned score, we substituted the derived prior score (19) into the process of Algorithm 2. The results indicate that Our Iter SDE CE & DD scheme outperforms traditional two-stage algorithms such as Pilot LS-CE & ZF-DD and Pilot LS-CE & OAMP-DD. However, it is less effective compared to the Iter LMMSE-CE & OAMP-DD algorithms. This is because, in a simple Rayleigh channel distribution, the mean and covariance matrix—representing first and second-order statistical information—are adequate to capture the channel’s distribution characteristics. Consequently, the LMMSE-CE and OAMP-DD achieve near-optimal performance, while the prior score offers limited additional information. Additionally, numerical errors in solving the SDE process can affect the scheme’s final performance. This suggests that the proposed scheme excels with more complex channel distributions, whereas existing methods already perform well in simple Rayleigh channels.
@@
Figure 12: NMSE and BER performance with respect to the pilot symbol length $L_{p}$ for different CE and DD algorithms.
Fig. 12 compares the performance curves of CE and DD as the functions of pilot length $L_{p}$ for different schemes. In the simulation, each user’s data length is $L_{d}\!=\!50$, and the received SNR at the BS is set to 10 dB. As shown in Fig. 12(a), when $L_{p}\!>\!12$, Our Iter SDE CE & DD can achieve the performance of SDE-CE & Perfect-Data, whereas the LS algorithm, which relies solely on pilots, exhibits a notable performance gap relative to iterative CE and DD algorithms. This suggests that additional computation can potentially reduce communication pilot overhead. Also Iter LMMSE-CE & OAMP-DD and LS-CE & Perfect-Data exhibit similar performance, indicating that the estimated data is sufficiently robust to enhance CE. Fig. 12(b) shows that the BER performance of Our Iter SDE CE & DD is considerably better than that of Perfect-CSI & OAMP-DD, which relies on the perfect CSI. As expected, Perfect-CSI & SDE-DD attains the best performance.
@@
Figure 13: NMSE and BER performance with respect to the data symbol length $L_{d}$ for different CE and DD algorithms.
Fig. 13 depicts the performance curves of CE and DD across various schemes as the functions of data length $L_{d}$. In the simulation, the pilot length for each UE is fixed at $L_{p}\!=\!15$, with the received $\text{SNR}\!=\!10$ dB at the BS. As can be seen from Fig. 13(a), the iterative algorithm improves CE performance as data length increases. For $L_{d}\!\geq\!30$, the NMSE of Our Iter SDE CE & DD matches that of SDE-CE & Perfect-Data, offering an NMSE gain of 1 to 2 dB over Iter LMMSE-CE & OAMP-DD. As depicted in Fig. 13(b), increasing $L_{d}$ improves the accuracy of CE, thereby enhancing the BER performance of Iter LMMSE-CE & OAMP-DD and Our Iter SDE CE & DD. Moreover, when $L_{d}\!=\!70$, the BER of Iter LMMSE-CE & OAMP-DD approximates that of Perfect-CSI & OAMP-DD, while Our Iter SDE CE & DD surpasses Perfect-CSI & OAMP-DD, in terms of BER.
@@
Figure 14: NMSE and BER performance with respect to the step numbers $N_{H}$ and $N_{X}$ of our Iter SDE CE & DD algorithm.
Fig. 14 compares the performance of CE and DD of Our Iter SDE CE & DD scheme as the function of the maximum algorithm iteration numbers $N_{H}$ and $N_{X}$. In the simulation, each UE’s pilot length is set to $L_{p}\!=\!15$ and the data length is $L_{d}\!=\!50$, with the received SNR fixed at 10 dB. For the predefined noise diffusion functions $\sigma(t)$ and $\tau(t)$, the interval $t\!\in\![0,\,1]$ can be divided into segments with different values of $N_{X}$ and $N_{H}$ to control the noise variance gap between adjacent steps, thereby balancing generation speed and performance. As shown in Fig. 14(a), as the channel sampling steps $N_{H}$ increase, the NMSE performance improves gradually due to a more elaborate denoising process, resulting in more accurate SDE solutions. When $N_{H}\!=\!1000$, the NMSE of Our Iter SDE CE & DD approximates that of SDE-CE & Perfect-Data. Moreover, under the same $N_{H}$, CE performance remains similar for different $N_{D}$, possibly due to SDE-based CE not being sensitive to DD errors. Additionally, Fig. 14(b) shows that as the data sampling steps $N_{D}$ increase, the BER performance of DD gradually improves. However, the BER curves vary for different $N_{H}$, indicating that for DD problems, SDE-based DD methods are more sensitive to CE errors.
@@
Fig. 15(a) illustrates the NMSE performance curve of Our Iter SDE CE & DD relative to the iteration index $i$ with varying pilot lengths $L_{p}$. The maximum number of iterations is set to $N_{H}=1500$ ($i\leq 1500$). The simulation results indicate that as the pilot length $L_{p}$ increases, the convergence iteration count for the proposed algorithm decreases from approximately 400 iterations at $L_{p}=12$ to about 100 iterations at $L_{p}=27$. Additionally, the final converged NMSE improves from 17.1-17.1 dB to 18.5-18.5 dB. Likewise, Fig. 15(b) demonstrates that with different SNRs, both the convergence speed and the final DD outcome of our algorithm improve as the SNR increases. Despite employing a maximum of 1500 iterations, the simulation results reveal that under favorable conditions, such as $L_{p}=27$ and $\text{SNR}=16$ dB, the proposed scheme requires only about 50 to 100 iterations to achieve satisfactory results.
@@
In Tab. 2, we present the average computation latency for one single realization of Our Iter SDE CE & DD algorithm, both with and without LMMSE initialization, across various pilot lengths. Simulations were performed using an NVIDIA GeForce GTX 4090 GPU and an AMD EPYC 7J13 CPU. As the pilot length $L_{p}$ increases, the computational latency of the method with LMMSE initialization gradually decreases, whereas the latency without LMMSE initialization remains relatively constant. This occurs because a longer pilot results in more accurate initial channel estimation, which corresponds to a smaller diffusion noise-level in the iteration process, resulting in fewer required iteration steps. Similarly, Tab. 3 illustrates the differences in computational latency across various SNRs when LMMSE initialization is applied or not. Notably, our approach requires only 29%-53% of the original scheme’s complexity, underscoring the effectiveness of the LMMSE initialization strategy.
Figure 16: Computation latency of different CE and DD schemes.
Fig. 16 compares the computational latency of various methods for processing a batch size of 64 in CE and DD tasks. Due to its asynchronous update mechanism and LMMSE initialization, Our Iter SDE CE & DD algorithm demonstrates significantly lower computational latency compared to Iter Langevin CE & DD algorithm. Furthermore, the latency of Our Iter SDE CE & DD is comparable to that of Iter LMMSE-CE & OAMP-DD, while the previous simulations have shown it outperforming Iter LMMSE-CE & OAMP-DD in both CE and DD. Although the latency of iterative CE and DD algorithms is higher than that of two-stage algorithms, such as Pilot LS-CE & ZF-DD and Pilot LS-CE & OAMP-DD, existing researches indicate that diffusion-based algorithms can reduce the number of steps from thousands to tens, or even to as few as 1 to 2 steps, through acceleration methods such as higher-order algorithms [48] and consistency models [50]. Therefore, accelerating diffusion algorithms still holds significant potential for addressing CE and DD problems.
@@ VII Conclusions
We have proposed techniques for addressing the challenges of AUD and JCEDD in massive random access, where standard deep learning often struggles to generalize across various problem dimensions using a single model. Our VPL-AUDNet architecture has been designed for variable pilot length transmission, enabling a single model to adapt to different pilot lengths and antennas, thereby achieving superior AUD performance in both 3GPP and Rayleigh channels. Furthermore, we have introduced a generative diffusion models-driven JCEDD scheme using PC samplers, which significantly enhances CE and DD performance through an asynchronous iterative algorithm. These proposed schemes have demonstrated the potential of utilizing deep learning and generative models to address signal detection and estimation challenges in massive random access. A future research direction of interest is to consider accelerating generative model methods to further reduce inference latency.

### v3

@@ VI-A4 Comparative Schemes
For AUD, the proposed method is compared with two baseline algorithms: AMP [13] and OMP [12]. Additionally, we also assess the performance of the VPL-AUDNet scheme trained with a fixed pilot length. For CE and $D$, the following algorithms are included in the comparisons:
@@
$\bullet$ Perfect-CSI & ZF-$D$: The CSI $\bm{H}_{a}$ is perfectly known, and employ ZF for $D$;
$\bullet$ Perfect-CSI & OAMP-$D$: The CSI $\bm{H}_{a}$ is perfectly known, and utilize OAMP [21] for $D$;
$\bullet$ Pilot LS-CE & ZF-$D$: Utilize pilots to perform LS CE, and use the estimated channel for ZF $D$;
$\bullet$ Pilot LS-CE & OAMP-$D$: Utilize pilots to perform LS CE, and use the estimated channel for OAMP [21] $D$;
$\bullet$ Iter LMMSE-CE & OAMP-$D$: Employ LMMSE for CE and OAMP for $D$, with CE and $D$ iteratively executed to enhance the performance of each other [21];
$\bullet$ Iter Langevin CE & $D$: Conduct iterative CE and $D$ using ALD [33];
@@
$\bullet$ Perfect-CSI & SDE-$D$: The CSI $\bm{H}_{a}$ is known, and apply SDE solver for $D$;
$\bullet$ Our Iter SDE CE & $D$: The proposed Algorithm 2.
@@ VI-C Performance of JCEDD
Figure 10: NMSE and BER performance with respect to the received SNR for different CE and $D$ algorithms under 3GPP channels.
@@
Fig. 10 illustrates the performance curves of CE and $D$ under varying SNRs for different schemes. In the simulations, each user is allocated a pilot length of $L_{p}\!=\!15$ and a data length of $L_{d}\!=\!50$. As shown in Fig. 10(a), the iterative CE and $D$ schemes (i.e., Iter Langevin CE & $D$, Iter LMMSE-CE & OAMP-$D$, and Our Iter SDE CE & $D$) achieve an NMSE gain of approximately 7 and 10 dB when compared to the non-iterative Pilot LS-CE & ZF-$D$, and Pilot LS-CE & OAMP-$D$ schemes. This indicates a significant gain in CE performance by using detected data to refine CE. As SNR increases, Iter LMMSE-CE & OAMP-$D$’s performance approaches that of LS-CE & Perfect-Data. Similarly, the performance of Our Iter SDE CE & $D$ (Algorithm 2) matches SDE-CE & Perfect-Data (Algorithm 1) for $\text{SNR}\!\geq\!8$ dB. When the data is not perfectly known, Our Iter SDE CE & $D$ outperforms Iter LMMSE-CE & OAMP-$D$ because it more effectively utilizes the prior distribution of the channel through a learned prior score function, whereas Iter LMMSE-CE & OAMP-$D$ relies only on the channel’s second-order statistical information via LMMSE. Additionally, Our Iter SDE CE & $D$ implements a PC sampler to enhance SDE solution performance beyond what Iter Langevin CE & $D$ algorithm can achieve with only a corrector.
@@
Fig. 10(b) compares the BER performance of various algorithms. Similar to Fig. 10(a), the non-iterative Pilot LS-CE & ZF-$D$ and Pilot LS-CE & OAMP-$D$ algorithms perform worse than the iterative algorithms. Our Iter SDE CE & $D$ scheme achieves performance close to that of Perfect-CSI & SDE-$D$ and surpasses the Iter LMMSE-CE & OAMP-$D$ and Iter Langevin CE & $D$ algorithms. At a BER of $10^{-4}$, Our Iter SDE CE & $D$ offers a 3 dB SNR gain over Iter LMMSE-CE & OAMP-$D$. Furthermore, Perfect-CSI & SDE-$D$ outperforms Perfect-CSI & OAMP-$D$ and Perfect-CSI & ZF-$D$ algorithms. This is likely because the channel matrix elements for the 3GPP scenario do not meet the i.i.d. distribution assumptions for $D$ tasks, causing OAMP algorithms to underperform. On the other hand, the ZF algorithm does not exploit the prior distribution information of the constellation. In contrast, Perfect-CSI & SDE-$D$ thoroughly leverages the prior distribution of the data and the likelihood information of the received signals, resulting in superior performance.
@@
Figure 11: NMSE and BER performance with respect to the received SNR for different CE and $D$ algorithms under the MIMO Rayleigh channel.
@@
In Fig. 11, we conducted simulations using the MIMO Rayleigh channel model, where the channel prior is assumed to follow $\mathcal{N}(\bm{0},\bm{I})$. Instead of utilizing a neural network-learned score, we substituted the derived prior score (19) into the process of Algorithm 2. The results indicate that Our Iter SDE CE & $D$ scheme outperforms traditional two-stage algorithms such as Pilot LS-CE & ZF-$D$ and Pilot LS-CE & OAMP-$D$. However, it is less effective compared to the Iter LMMSE-CE & OAMP-$D$ algorithms. This is because, in a simple Rayleigh channel distribution, the mean and covariance matrix—representing first and second-order statistical information—are adequate to capture the channel’s distribution characteristics. Consequently, the LMMSE-CE and OAMP-$D$ achieve near-optimal performance, while the prior score offers limited additional information. Additionally, numerical errors in solving the SDE process can affect the scheme’s final performance. This suggests that the proposed scheme excels with more complex channel distributions, whereas existing methods already perform well in simple Rayleigh channels.
@@
Figure 12: NMSE and BER performance with respect to the pilot symbol length $L_{p}$ for different CE and $D$ algorithms.
Fig. 12 compares the performance curves of CE and $D$ as the functions of pilot length $L_{p}$ for different schemes. In the simulation, each user’s data length is $L_{d}\!=\!50$, and the received SNR at the BS is set to 10 dB. As shown in Fig. 12(a), when $L_{p}\!>\!12$, Our Iter SDE CE & $D$ can achieve the performance of SDE-CE & Perfect-Data, whereas the LS algorithm, which relies solely on pilots, exhibits a notable performance gap relative to iterative CE and $D$ algorithms. This suggests that additional computation can potentially reduce communication pilot overhead. Also Iter LMMSE-CE & OAMP-$D$ and LS-CE & Perfect-Data exhibit similar performance, indicating that the estimated data is sufficiently robust to enhance CE. Fig. 12(b) shows that the BER performance of Our Iter SDE CE & $D$ is considerably better than that of Perfect-CSI & OAMP-$D$, which relies on the perfect CSI. As expected, Perfect-CSI & SDE-$D$ attains the best performance.
@@
Figure 13: NMSE and BER performance with respect to the data symbol length $L_{d}$ for different CE and $D$ algorithms.
Fig. 13 depicts the performance curves of CE and $D$ across various schemes as the functions of data length $L_{d}$. In the simulation, the pilot length for each UE is fixed at $L_{p}\!=\!15$, with the received $\text{SNR}\!=\!10$ dB at the BS. As can be seen from Fig. 13(a), the iterative algorithm improves CE performance as data length increases. For $L_{d}\!\geq\!30$, the NMSE of Our Iter SDE CE & $D$ matches that of SDE-CE & Perfect-Data, offering an NMSE gain of 1 to 2 dB over Iter LMMSE-CE & OAMP-$D$. As depicted in Fig. 13(b), increasing $L_{d}$ improves the accuracy of CE, thereby enhancing the BER performance of Iter LMMSE-CE & OAMP-$D$ and Our Iter SDE CE & $D$. Moreover, when $L_{d}\!=\!70$, the BER of Iter LMMSE-CE & OAMP-$D$ approximates that of Perfect-CSI & OAMP-$D$, while Our Iter SDE CE & $D$ surpasses Perfect-CSI & OAMP-$D$, in terms of BER.
@@
Figure 14: NMSE and BER performance with respect to the step numbers $N_{H}$ and $N_{X}$ of our Iter SDE CE & $D$ algorithm.
Fig. 14 compares the performance of CE and $D$ of Our Iter SDE CE & $D$ scheme as the function of the maximum algorithm iteration numbers $N_{H}$ and $N_{X}$. In the simulation, each UE’s pilot length is set to $L_{p}\!=\!15$ and the data length is $L_{d}\!=\!50$, with the received SNR fixed at 10 dB. For the predefined noise diffusion functions $\sigma(t)$ and $\tau(t)$, the interval $t\!\in\![0,\,1]$ can be divided into segments with different values of $N_{X}$ and $N_{H}$ to control the noise variance gap between adjacent steps, thereby balancing generation speed and performance. As shown in Fig. 14(a), as the channel sampling steps $N_{H}$ increase, the NMSE performance improves gradually due to a more elaborate denoising process, resulting in more accurate SDE solutions. When $N_{H}\!=\!1000$, the NMSE of Our Iter SDE CE & $D$ approximates that of SDE-CE & Perfect-Data. Moreover, under the same $N_{H}$, CE performance remains similar for different $N_{D}$, possibly due to SDE-based CE not being sensitive to $D$ errors. Additionally, Fig. 14(b) shows that as the data sampling steps $N_{D}$ increase, the BER performance of $D$ gradually improves. However, the BER curves vary for different $N_{H}$, indicating that for $D$ problems, SDE-based $D$ methods are more sensitive to CE errors.
@@
Fig. 15(a) illustrates the NMSE performance curve of Our Iter SDE CE & $D$ relative to the iteration index $i$ with varying pilot lengths $L_{p}$. The maximum number of iterations is set to $N_{H}=1500$ ($i\leq 1500$). The simulation results indicate that as the pilot length $L_{p}$ increases, the convergence iteration count for the proposed algorithm decreases from approximately 400 iterations at $L_{p}=12$ to about 100 iterations at $L_{p}=27$. Additionally, the final converged NMSE improves from 17.1-17.1 dB to 18.5-18.5 dB. Likewise, Fig. 15(b) demonstrates that with different SNRs, both the convergence speed and the final $D$ outcome of our algorithm improve as the SNR increases. Despite employing a maximum of 1500 iterations, the simulation results reveal that under favorable conditions, such as $L_{p}=27$ and $\text{SNR}=16$ dB, the proposed scheme requires only about 50 to 100 iterations to achieve satisfactory results.
@@
In Tab. 2, we present the average computation latency for one single realization of Our Iter SDE CE & $D$ algorithm, both with and without LMMSE initialization, across various pilot lengths. Simulations were performed using an NVIDIA GeForce GTX 4090 GPU and an AMD EPYC 7J13 CPU. As the pilot length $L_{p}$ increases, the computational latency of the method with LMMSE initialization gradually decreases, whereas the latency without LMMSE initialization remains relatively constant. This occurs because a longer pilot results in more accurate initial channel estimation, which corresponds to a smaller diffusion noise-level in the iteration process, resulting in fewer required iteration steps. Similarly, Tab. 3 illustrates the differences in computational latency across various SNRs when LMMSE initialization is applied or not. Notably, our approach requires only 29%-53% of the original scheme’s complexity, underscoring the effectiveness of the LMMSE initialization strategy.
Figure 16: Computation latency of different CE and $D$ schemes.
Fig. 16 compares the computational latency of various methods for processing a batch size of 64 in CE and $D$ tasks. Due to its asynchronous update mechanism and LMMSE initialization, Our Iter SDE CE & $D$ algorithm demonstrates significantly lower computational latency compared to Iter Langevin CE & $D$ algorithm. Furthermore, the latency of Our Iter SDE CE & $D$ is comparable to that of Iter LMMSE-CE & OAMP-$D$, while the previous simulations have shown it outperforming Iter LMMSE-CE & OAMP-$D$ in both CE and $D$. Although the latency of iterative CE and $D$ algorithms is higher than that of two-stage algorithms, such as Pilot LS-CE & ZF-$D$ and Pilot LS-CE & OAMP-$D$, existing researches indicate that diffusion-based algorithms can reduce the number of steps from thousands to tens, or even to as few as 1 to 2 steps, through acceleration methods such as higher-order algorithms [48] and consistency models [50]. Therefore, accelerating diffusion algorithms still holds significant potential for addressing CE and $D$ problems.
@@ VII Conclusions
We have proposed techniques for addressing the challenges of AUD and JCEDD in massive random access, where standard deep learning often struggles to generalize across various problem dimensions using a single model. Our VPL-AUDNet architecture has been designed for variable pilot length transmission, enabling a single model to adapt to different pilot lengths and antennas, thereby achieving superior AUD performance in both 3GPP and Rayleigh channels. Furthermore, we have introduced a generative diffusion models-driven JCEDD scheme using PC samplers, which significantly enhances CE and $D$ performance through an asynchronous iterative algorithm. These proposed schemes have demonstrated the potential of utilizing deep learning and generative models to address signal detection and estimation challenges in massive random access. A future research direction of interest is to consider accelerating generative model methods to further reduce inference latency.